# CHEMOUT: CHEMical complexity in star-forming regions of the OUTer Galaxy

### v2

@@ 4.1.1 Formaldehyde and methanol updated fits
The fits for H2CO and CH3OH have been revised due to an incorrect assumption in fontani2022chemoutII, which originated from a conversion between $T_{\mathrm{A}}^{*}$ and $T_{\rm MB}$. Since the beam efficiency $\eta_{\text{eff}}$ of the telescope was not explicitly specified in bernal2021methanol, the values were taken from Appendix C.3 of the ARO 12m user manual; however, outdated efficiencies were inadvertently used. Updated efficiencies [^3] From https://aro.as.arizona.edu/?q=beam-efficiencies. have now been adopted to estimate the $T_{\mathrm{MB}}$ of these two species. For H2CO at 2 mm we divided $T_{\mathrm{A}}^{*}$ by 0.81, and for CH3OH at 3 mm we divided by 0.91. The source sizes for H2CO and CH3OH were re-estimated using Eq. 3 in fontani2002structure. Consequently, the MADCUBA fits were recalculated, and the resulting parameters are reported in Appendix C. The excitation temperatures of CH3OH are listed in Table 5.
@@ 4.2 Fractional abundances
The plots showing $X_{\text{mol}}$ against $R_{\text{GC}}$ for each molecule are shown in Fig. 1. For comparison, the fractional abundances calculated using a gas-to-dust ratio $\gamma=100$ are also shown. A linear fit is performed for each plot, and the Pearson correlation coefficient [^5] This coefficient measures the strength and direction of the linear relationship between two variables, X and Y, and it is calculated as the covariance of X and Y divided by the product of their standard deviations. It ranges from 1-1 to +1+1, where +1+1 indicates perfect positive correlation, 0 indicates no linear correlation, and 1-1 indicates perfect negative correlation., $\rho$, is calculated to inspect possible (anti-)correlations. The linear fit applied to the data of fractional abundances, calculated using a gas-to-dust ratio from giannetti2017galactocentric, gives:
@@
Figure 1: Galactocentric gradients of fractional abundances with respect to H2, $X_{\text{mol}}$. The plot shows the trends for HCN (a), HCO+(b), c-C3H2(c), H13CO+(d), HCO (e), SO (f), H2CO (g), and CH3OH (h) as a function of the galactocentric radius ($R_{\text{GC}}$). The light blue data represent the abundances calculated using a constant gas-to-dust ratio ($\gamma=100$), while the dark blue data illustrate those estimated using a non-constant gas-to-dust ratio ($\gamma=\gamma(R_{\text{GC}})$ from giannetti2017galactocentric). For both datasets, the linear regression results are shown as the light blue and dark blue lines, respectively. The $1\sigma$ error bars over the slope of the gradients are displayed for each fit. The upper-limit values are represented with triangles. The gradients of the elemental abundances of carbon (C and 13C), nitrogen (N), oxygen (O), and sulfur (S), as reported by mendez2022gradients, are plotted as dashed lines. The product of the parent elements of the species is represented, in each subplot, by a dash–dotted black line. All elemental trends are plotted with a vertical offset to align with the starting point of the molecular gradients derived using $\gamma(R_{\text{GC}})$, to facilitate comparison of their slopes with the linear fit of the molecular abundances. In the upper-right side of each subplot, the Pearson correlation coefficient, $\rho$, is shown (estimated only for the abundances estimated using the non-constant gas-to-dust ratio).
@@ 5.1 Galactocentric gradients of fractional abundances
The trends obtained for the molecular abundance (with respect to H2) were compared with the galactocentric abundance gradients of carbon (both 12C and 13C), oxygen, nitrogen, and sulfur. These are constituting elements of the studied species, and the gradients were determined in mendez2022gradients through the analysis of the spectra of Galactic H II regions. The galactocentric gradient of 13C was estimated by coupling the carbon gradient with the galactocentric trend of the $^{13}\text{C}/^{12}\text{C}$ ratio from yan2023direct. Among 8.34 kpc, the galactocentric distance of the Sun, and 24 kpc, which is the distance of the most distant source in our sample, the abundance ratios [C/H], [13C/H], [N/H], [O/H], and [S/H] decrease by factors of approximately $\sim 16$, $\sim 36$, $\sim 10$, $\sim 5$, and $\sim 4$, respectively. Extrapolating the gradients derived in the OG (Eqs. 38) to the Solar circle, over the same range of $R_{\text{GC}}$, we find that the abundance of the studied species, i.e., HCN, HCO+, c-C3H2, H13CO+, HCO, SO, H2CO, and CH3OH decrease by a factor of $\sim 44$, $\sim 17$, $\sim 8$, $\sim 4$, $\sim 8$, $\sim 18$, $\sim 41$, and $\sim 194.$
@@
In contrast, the decreasing rate of SO appears steeper than that of [S/H]. For SO, the higher decreasing rate with respect to [S/H] indicates a lower efficiency of the formation of these molecules in the OG with respect to the local Galaxy. We speculate that a steeper decreasing trend for SO, which is the best shock tracer among those that we analyzed in this work, could indicate that, in the OG, outflows have a reduced influence on the gas composition compared to the local Galaxy. This result would be consistent with the decreasing line width with $R_{\text{GC}}$ (Sect. 4.3). Another shock tracer is CH3OH, which, although its decreasing trend is marginally consistent with that of carbon, exhibits the most pronounced reduction among the species investigated in the OG. The steep decreasing trend of CH3OH could also arise from a less efficient grain-surface chemistry in the OG. The results concerning HCO, H2CO, and CH3OH, obtained with the updated abundances (and distances), confirm the ones of fontani2022chemoutII. In fontani2024chemout, a correlation between the abundances of CH3OH and SO was found, which is confirmed over the entire CHEMOUT sample, although no correlation between the line widths of these two species is observed, hinting that they are not tracing the same gas components.
The decreasing rate of H13CO+, although lower than that of [13C/H], aligns well with the decreasing rate observed for oxygen. This correlation is not observed for HCO+, whose decreasing rate is consistent with carbon rather than oxygen. This discrepancy could suggest a possible revision of the galactocentric gradient of the 13C/12C ratio, with a shallower trend at high $R_{\text{GC}}$, in agreement with the Galactic chemical evolution discussed in colzi2022chemout. Moreover, the detailed analysis of the 13C/12C ratios from the HCO+ and H13CO+ lines will be the subject of a forthcoming paper (Colzi et al. in prep.). Therefore, the results regarding c-C3H2, HCO, and especially H13CO+ lead to the puzzling conclusion of observing more carbon-bearing molecules than the available carbon atoms in the OG, their decreasing rates being lower than that of [C/H] (or [13C/H]).
@@ 5.2 Galactocentric gradients of line widths
This result is also consistent with the study of lin2025inadequate, where the velocity dispersion of low-$J$ 13CO lines has been studied toward molecular clouds in the low-metallicity part of the Galactic disk. A decreasing trend of the line width is observed with $R_{\text{GC}}$, and the sub-virial dynamics observed in the metal-poor clouds hint at how the gas turbulence is not sufficient to counterbalance the cloud’s self-gravity. As stated in lin2025inadequate, in low-metallicity environments, the magnetic field may play a more dominant role, replacing gas turbulence in sustaining the clouds.
@@ 5.3 Galactocentric gradients of molecular ratios
Studies on the formation pathways of HCO+, H13CO+ (klemperer1970carrier; millar1991gas), HCO, H2CO, and CH3OH (watanabe2002efficient; fuchs2009hydrogenation; minissale2016dust; rivilla2019first; jimenez2025modelling) have shown that their formation is linked to the presence of CO. In contrast, c-C3H2 is formed from ”free” carbon (not bound in CO) (thaddeus1985laboratory; sipila2016understanding). Moreover, it is known that [C/O] decreases with distance from the Galactic center (mendez2022gradients). Therefore, we expect to have less free carbon in the OG to produce c-C3H2, because most of it would be locked in CO. This would lead to a prediction of an increasing trend for the CO/c-C3H2 ratio.
@@ 5.4 Caveats and uncertainties
The comparison between observations and chemical models by fontani2024chemout for WB89-670 suggests that the methanol could be sub-thermally excited: the kinetic temperatures predicted by chemical modeling are in the $\sim 20-40\penalty 10000\$K range, while the excitation temperatures are on the order of $\sim 10$ K. Therefore, an important caveat arises from the assumed $T_{\text{ex}}$ for the species for which we cannot estimate it directly from the observations.
Furthermore, the sources are located in a range of heliocentric distances, $d$, between 3 and 15.1 kpc, which implies that the linear scale probed is not the same in all targets. fontani2024chemout demonstrated the presence of a high differentiation at smaller angular scales in WB89-670, and hence the average column densities derived in this work refer to linear scales with (eventually) different average conditions. However, this caveat is typical of source surveys, and in any case, we compare emissions arising from core envelopes in all targets.
Some results may also be influenced by low statistics, particularly within certain $R_{\text{GC}}$ ranges, where the data characterizing a few sources have a bigger impact on the gradients. This concern is more significant for formaldehyde and methanol, given their even lower sampling rates (only 18 molecular clumps of 39 are observed in H2CO and CH3OH spectral windows), over a smaller $R_{\text{GC}}$ range (fontani2022chemoutII).
@@
Moreover, in the comparison between molecular and elemental trends, a further caveat comes from the fact that the elemental abundance gradients are extrapolated in the extreme OG ($R_{\text{GC}}\gtrsim 16$ kpc), where observations are sparse and thus the gradients are poorly constrained, and could be different from those at $R_{\text{GC}}\leq 16$ kpc. Indeed, recent studies have reported evidence of a flattening of the metallicity gradient at large galactocentric radii, which could influence the observed molecular abundances and potentially explain the unexpectedly high chemical complexity in the outskirts of the Galaxy (donor2020open; myers2022open; magrini2023gaia; da2023oxygen).
@@ 6 Conclusions and outlooks
In this work, we studied the chemical composition of 35 star-forming regions in the OG ($9\leq R_{\text{GC}}\leq 24$ kpc), using IRAM 30 m telescope data from the CHEMOUT project. We identified 39 molecular clumps and analyzed the molecular species detected in the greatest number of targets across our sample; i.e., HCN (39/39), HCO+ (39/39), c-C3H2 (38/39), H13CO+ (30/39), HCO (29/39), and SO (25/39). We estimated their column densities and calculated the fractional abundances of these species with respect to H2. We also updated the abundances of H2CO (18/18) and CH3OH (17/18) from fontani2022chemoutII, using the H2 column densities obtained from the new NIKA2 continuum maps.
The fractional abundances of the studied species with respect to H2 estimated assuming a gas-to-dust ratio dependent on $R_{\text{GC}}$ (giannetti2017galactocentric) show a decreasing trend with the galactocentric radius. For HCN, HCO+, c-C3H2, HCO, SO, H2CO, and CH3OH the abundances scale as the elemental abundance of carbon ([C/H]) between the solar circle and $\sim 24\penalty 10000\$kpc. This indicates that the efficiency in the formation of these molecules, which scales with the availability of the parent element (i.e., C), is at least as high as that found in the local Galaxy. Instead, the decreasing rate of SO is higher than that of [S/H], indicating a lower efficiency in the formation of this species in the OG. Conversely, the decreasing rate of H13CO+ presents the puzzling result that it is lower than that of [13C/H], resulting in a higher formation efficiency in the OG compared to the inner Galaxy.
@@
The next step for the CHEMOUT project is to infer a more precise kinetic temperature of the sources to better estimate the column densities from the line analysis. Determining this parameter will also help to reduce the space parameters on a chemical modeling analysis (e.g., using UCLCHEM; holdship2017uclchem), which is crucial to inferring the environmental differences of the star-forming regions in the different areas of the Galaxy and understanding the parameters driving the chemistry (vermarien2025understanding). The comparison between high-resolution observations and chemical models is crucial to disentangling the several contributions to these formation processes.
@@ Appendix A The CHEMOUT sample
Source aaaaaaThe sources labeled with a ”-b” are related to secondary velocity features detected in the spectra, likely due to a second molecular clump within the telescope beam. R.A. Dec $\ell$ $V_{\text{LSR}}$bbbbbbLocal standard of rest velocities estimated from the velocity peak of the c-C3H2 $J_{K_{a},K_{c}}=2_{1,2}-1_{0,1}$ line, from fontani2022chemout. For the source WB89-315, being undetected in c-C3H2, we use the $V_{\text{LSR}}$ given in blair2008formaldehyde. $R_{\text{GC}}$ccccccGalactocentric distances and heliocentric distances, respectively, based on the rotation curve of russeil2017milky, and the $V_{\text{LSR}}$ values shown in this table. For WB89-437 and WB89-621, we assume the more accurate distances estimated through maser parallax measurements in hachisuka2015parallaxes. Their kinematic galactocentric distances are 15.7 and 18.9 kpc, respectively. ddccccccGalactocentric distances and heliocentric distances, respectively, based on the rotation curve of russeil2017milky, and the $V_{\text{LSR}}$ values shown in this table. For WB89-437 and WB89-621, we assume the more accurate distances estimated through maser parallax measurements in hachisuka2015parallaxes. Their kinematic galactocentric distances are 15.7 and 18.9 kpc, respectively. $N_{\text{H}_{2}}$ddddddDerived from the NIKA2 continuum maps as described in Sect. 3.2. $T_{\text{ex}}$eeeeeeExcitation temperature of CH3OH, derived from the updated fits of the data presented in fontani2022chemoutII (see Sect. 4.1.1), adopted to fit the other species.

### v3

@@ 4.1.1 Formaldehyde and methanol updated fits
The fits for H2CO and CH3OH have been revised due to an incorrect assumption in Fontani et al. (2022b), which originated from a conversion between $T_{\mathrm{A}}^{*}$ and $T_{\rm MB}$. Since the beam efficiency $\eta_{\text{eff}}$ of the telescope was not explicitly specified in Bernal et al. (2021), the values were taken from Appendix C.3 of the ARO 12m user manual; however, outdated efficiencies were inadvertently used. Updated efficiencies [^3] From https://aro.as.arizona.edu/?q=beam-efficiencies. have now been adopted to estimate the $T_{\mathrm{MB}}$ of these two species. For H2CO at 2 mm we divided $T_{\mathrm{A}}^{*}$ by 0.81, and for CH3OH at 3 mm we divided by 0.91. The source sizes for H2CO and CH3OH were re-estimated using Eq. 3 in Fontani et al. (2002). Consequently, the MADCUBA fits were recalculated, and the resulting parameters are reported in Appendix C. The excitation temperatures of CH3OH are listed in Table 5.
@@ 4.2 Fractional abundances
The plots showing $X_{\text{mol}}$ against $R_{\text{GC}}$ for each molecule are shown in Fig. 1. For comparison, the fractional abundances calculated using a gas-to-dust ratio $\gamma=100$ are also shown. A linear fit is performed for each plot, and the Pearson correlation coefficient [^5] This coefficient measures the strength and direction of the linear relationship between two variables, X and Y, and it is calculated as the covariance of X and Y divided by the product of their standard deviations. It ranges from 1-1 to +1+1, where +1+1 indicates perfect positive correlation, 0 indicates no linear correlation, and 1-1 indicates perfect negative correlation., $\rho$, is calculated to inspect possible (anti-)correlations. The linear fit applied to the data of fractional abundances, calculated using a gas-to-dust ratio from Giannetti et al. (2017), gives:
@@
Figure 1: Galactocentric gradients of fractional abundances with respect to H2, $X_{\text{mol}}$. The plot shows the trends for HCN (a), HCO+(b), c-C3H2(c), H13CO+(d), HCO (e), SO (f), H2CO (g), and CH3OH (h) as a function of the galactocentric radius ($R_{\text{GC}}$). The light blue data represent the abundances calculated using a constant gas-to-dust ratio ($\gamma=100$), while the dark blue data illustrate those estimated using a non-constant gas-to-dust ratio ($\gamma=\gamma(R_{\text{GC}})$ from Giannetti et al. 2017). For both datasets, the linear regression results are shown as the light blue and dark blue lines, respectively. The $1\sigma$ error bars over the slope of the gradients are displayed for each fit. The upper-limit values are represented with triangles. The gradients of the elemental abundances of carbon (C and 13C), nitrogen (N), oxygen (O), and sulfur (S), as reported by Méndez-Delgado et al. (2022), are plotted as dashed lines. The product of the parent elements of the species is represented, in each subplot, by a dash–dotted black line. All elemental trends are plotted with a vertical offset to align with the starting point of the molecular gradients derived using $\gamma(R_{\text{GC}})$, to facilitate comparison of their slopes with the linear fit of the molecular abundances. In the upper-right side of each subplot, the Pearson correlation coefficient, $\rho$, is shown (estimated only for the abundances estimated using the non-constant gas-to-dust ratio).
@@ 5.1 Galactocentric gradients of fractional abundances
The trends obtained for the molecular abundance (with respect to H2) were compared with the galactocentric abundance gradients of carbon (both 12C and 13C), oxygen, nitrogen, and sulfur. These are constituting elements of the studied species, and the gradients were determined in Méndez-Delgado et al. (2022) through the analysis of the spectra of Galactic H II regions. The galactocentric gradient of 13C was estimated by coupling the carbon gradient with the galactocentric trend of the $^{13}\text{C}/^{12}\text{C}$ ratio from Yan et al. (2023). Among 8.34 kpc, the galactocentric distance of the Sun, and 24 kpc, which is the distance of the most distant source in our sample, the abundance ratios [C/H], [13C/H], [N/H], [O/H], and [S/H] decrease by factors of approximately $\sim 16$, $\sim 36$, $\sim 10$, $\sim 5$, and $\sim 4$, respectively. Extrapolating the gradients derived in the OG (Eqs. 38) to the Solar circle, over the same range of $R_{\text{GC}}$, we find that the abundance of the studied species, i.e., HCN, HCO+, c-C3H2, H13CO+, HCO, SO, H2CO, and CH3OH decrease by a factor of $\sim 44$, $\sim 17$, $\sim 8$, $\sim 4$, $\sim 8$, $\sim 18$, $\sim 41$, and $\sim 194.$
@@
In contrast, the decreasing rate of SO appears steeper than that of [S/H]. For SO, the higher decreasing rate with respect to [S/H] indicates a lower efficiency of the formation of these molecules in the OG with respect to the local Galaxy. We speculate that a steeper decreasing trend for SO, which is the best shock tracer among those that we analyzed in this work, could indicate that, in the OG, outflows have a reduced influence on the gas composition compared to the local Galaxy. This result would be consistent with the decreasing line width with $R_{\text{GC}}$ (Sect. 4.3). Another shock tracer is CH3OH, which, although its decreasing trend is marginally consistent with that of carbon, exhibits the most pronounced reduction among the species investigated in the OG. The steep decreasing trend of CH3OH could also arise from a less efficient grain-surface chemistry in the OG. The results concerning HCO, H2CO, and CH3OH, obtained with the updated abundances (and distances), confirm the ones of Fontani et al. (2022b). In Fontani et al. (2024), a correlation between the abundances of CH3OH and SO was found, which is confirmed over the entire CHEMOUT sample, although no correlation between the line widths of these two species is observed, hinting that they are not tracing the same gas components.
The decreasing rate of H13CO+, although lower than that of [13C/H], aligns well with the decreasing rate observed for oxygen. This correlation is not observed for HCO+, whose decreasing rate is consistent with carbon rather than oxygen. This discrepancy could suggest a possible revision of the galactocentric gradient of the 13C/12C ratio, with a shallower trend at high $R_{\text{GC}}$, in agreement with the Galactic chemical evolution discussed in Colzi et al. (2022). Moreover, the detailed analysis of the 13C/12C ratios from the HCO+ and H13CO+ lines will be the subject of a forthcoming paper (Colzi et al. in prep.). Therefore, the results regarding c-C3H2, HCO, and especially H13CO+ lead to the puzzling conclusion of observing more carbon-bearing molecules than the available carbon atoms in the OG, their decreasing rates being lower than that of [C/H] (or [13C/H]).
@@ 5.2 Galactocentric gradients of line widths
This result is also consistent with the study of Lin et al. (2025), where the velocity dispersion of low-$J$ 13CO lines has been studied toward molecular clouds in the low-metallicity part of the Galactic disk. A decreasing trend of the line width is observed with $R_{\text{GC}}$, and the sub-virial dynamics observed in the metal-poor clouds hint at how the gas turbulence is not sufficient to counterbalance the cloud’s self-gravity. As stated in Lin et al. (2025), in low-metallicity environments, the magnetic field may play a more dominant role, replacing gas turbulence in sustaining the clouds.
@@ 5.3 Galactocentric gradients of molecular ratios
Studies on the formation pathways of HCO+, H13CO+ (Klemperer 1970; Millar et al. 1991), HCO, H2CO, and CH3OH (Watanabe and Kouchi 2002; Fuchs et al. 2009; Minissale et al. 2016; Rivilla et al. 2019; Jiménez-Serra et al. 2025) have shown that their formation is linked to the presence of CO. In contrast, c-C3H2 is formed from ”free” carbon (not bound in CO) (Thaddeus et al. 1985; Sipilä et al. 2016). Moreover, it is known that [C/O] decreases with distance from the Galactic center (Méndez-Delgado et al. 2022). Therefore, we expect to have less free carbon in the OG to produce c-C3H2, because most of it would be locked in CO. This would lead to a prediction of an increasing trend for the CO/c-C3H2 ratio.
@@ 5.4 Caveats and uncertainties
The comparison between observations and chemical models by Fontani et al. (2024) for WB89-670 suggests that the methanol could be sub-thermally excited: the kinetic temperatures predicted by chemical modeling are in the $\sim 20-40\penalty 10000\$K range, while the excitation temperatures are on the order of $\sim 10$ K. Therefore, an important caveat arises from the assumed $T_{\text{ex}}$ for the species for which we cannot estimate it directly from the observations.
Furthermore, the sources are located in a range of heliocentric distances, $d$, between 3 and 15.1 kpc, which implies that the linear scale probed is not the same in all targets. Fontani et al. (2024) demonstrated the presence of a high differentiation at smaller angular scales in WB89-670, and hence the average column densities derived in this work refer to linear scales with (eventually) different average conditions. However, this caveat is typical of source surveys, and in any case, we compare emissions arising from core envelopes in all targets.
Some results may also be influenced by low statistics, particularly within certain $R_{\text{GC}}$ ranges, where the data characterizing a few sources have a bigger impact on the gradients. This concern is more significant for formaldehyde and methanol, given their even lower sampling rates (only 18 molecular clumps of 39 are observed in H2CO and CH3OH spectral windows), over a smaller $R_{\text{GC}}$ range (Fontani et al. 2022b).
@@
Moreover, in the comparison between molecular and elemental trends, a further caveat comes from the fact that the elemental abundance gradients are extrapolated in the extreme OG ($R_{\text{GC}}\gtrsim 16$ kpc), where observations are sparse and thus the gradients are poorly constrained, and could be different from those at $R_{\text{GC}}\leq 16$ kpc. Indeed, recent studies have reported evidence of a flattening of the metallicity gradient at large galactocentric radii, which could influence the observed molecular abundances and potentially explain the unexpectedly high chemical complexity in the outskirts of the Galaxy (Donor et al. 2020; Myers et al. 2022; Magrini et al. 2023; Da Silva et al. 2023).
@@ 6 Conclusions and outlooks
In this work, we studied the chemical composition of 35 star-forming regions in the OG ($9\leq R_{\text{GC}}\leq 24$ kpc), using IRAM 30 m telescope data from the CHEMOUT project. We identified 39 molecular clumps and analyzed the molecular species detected in the greatest number of targets across our sample; i.e., HCN (39/39), HCO+ (39/39), c-C3H2 (38/39), H13CO+ (30/39), HCO (29/39), and SO (25/39). We estimated their column densities and calculated the fractional abundances of these species with respect to H2. We also updated the abundances of H2CO (18/18) and CH3OH (17/18) from Fontani et al. (2022b), using the H2 column densities obtained from the new NIKA2 continuum maps.
The fractional abundances of the studied species with respect to H2 estimated assuming a gas-to-dust ratio dependent on $R_{\text{GC}}$ (Giannetti et al. 2017) show a decreasing trend with the galactocentric radius. For HCN, HCO+, c-C3H2, HCO, SO, H2CO, and CH3OH the abundances scale as the elemental abundance of carbon ([C/H]) between the solar circle and $\sim 24\penalty 10000\$kpc. This indicates that the efficiency in the formation of these molecules, which scales with the availability of the parent element (i.e., C), is at least as high as that found in the local Galaxy. Instead, the decreasing rate of SO is higher than that of [S/H], indicating a lower efficiency in the formation of this species in the OG. Conversely, the decreasing rate of H13CO+ presents the puzzling result that it is lower than that of [13C/H], resulting in a higher formation efficiency in the OG compared to the inner Galaxy.
@@
The next step for the CHEMOUT project is to infer a more precise kinetic temperature of the sources to better estimate the column densities from the line analysis. Determining this parameter will also help to reduce the space parameters on a chemical modeling analysis (e.g., using UCLCHEM; Holdship et al. 2017), which is crucial to inferring the environmental differences of the star-forming regions in the different areas of the Galaxy and understanding the parameters driving the chemistry (Vermariën et al. 2025). The comparison between high-resolution observations and chemical models is crucial to disentangling the several contributions to these formation processes.
@@ Appendix A The CHEMOUT sample
Source aaaaaaThe sources labeled with a ”-b” are related to secondary velocity features detected in the spectra, likely due to a second molecular clump within the telescope beam. R.A. Dec $\ell$ $V_{\text{LSR}}$bbbbbbLocal standard of rest velocities estimated from the velocity peak of the c-C3H2 $J_{K_{a},K_{c}}=2_{1,2}-1_{0,1}$ line, from Fontani et al. (2022a). For the source WB89-315, being undetected in c-C3H2, we use the $V_{\text{LSR}}$ given in Blair et al. (2008). $R_{\text{GC}}$ccccccGalactocentric distances and heliocentric distances, respectively, based on the rotation curve of Russeil et al. (2017), and the $V_{\text{LSR}}$ values shown in this table. For WB89-437 and WB89-621, we assume the more accurate distances estimated through maser parallax measurements in Hachisuka et al. (2015). Their kinematic galactocentric distances are 15.7 and 18.9 kpc, respectively. ddccccccGalactocentric distances and heliocentric distances, respectively, based on the rotation curve of Russeil et al. (2017), and the $V_{\text{LSR}}$ values shown in this table. For WB89-437 and WB89-621, we assume the more accurate distances estimated through maser parallax measurements in Hachisuka et al. (2015). Their kinematic galactocentric distances are 15.7 and 18.9 kpc, respectively. $N_{\text{H}_{2}}$ddddddDerived from the NIKA2 continuum maps as described in Sect. 3.2. $T_{\text{ex}}$eeeeeeExcitation temperature of CH3OH, derived from the updated fits of the data presented in Fontani et al. (2022b) (see Sect. 4.1.1), adopted to fit the other species.